RyeBread1:

The first step when dividing 9m^2 − 4m − 6 by 3m is shown. Which could be the next step? 3m − 3/4 −m/2 3m − 3/4 − 2/m 3m − 4/3 − m/2 3m − 4/3 − 2/m

RyeBread1:

I think the answer is B.

TheRobin:

I don't know

TheRobin:

but I'll make it even between me and @darkknight

RyeBread1:

ha thanks

RyeBread1:

@justjm

justjm:

I'll help you in a bit, or perhaps you can ask @Hero

RyeBread1:

i already did ask him he didnt respond

RyeBread1:

@Hero

justjm:

@InsatiableSuffering can help you out too

RyeBread1:

Ok.

InsatiableSuffering:

So we can treat this as a fraction. By dividing the quadratic by 3m, we would find that our problem would look like this\[\frac{ 9m ^{2}-4m-6 }{ 3m }\] which is equivalent to \[\frac{ 9m ^{2} }{3m}-\frac{ 4m }{ 3m }-\frac{ 6 }{ 3m }\] By solving it out, you would find that you would get \[3m-\frac{ 4 }{ 3 }-\frac{ 2 }{ m }\] All it takes it cancelling out terms that can be simplified and you have your answer.

RyeBread1:

thank you